OpenStudy (anonymous):

Evaluate the integral using integration by parts where possible. (X+1)e^x dx from 0 to 1

OpenStudy (jhannybean):

\[\large \int_{0}^{1} (x+1)\cdot e^{x}dx\]

OpenStudy (jhannybean):

\[\large {u = (x+1)\\du=dx}\]\[\large {dv=e^xdx\\v=e^x}\]

OpenStudy (jhannybean):

Can you solve it from here?

OpenStudy (dan815):



OpenStudy (jhannybean):

\[\large e^{x}(x+1)]_{0}^{1} - \int_{0}^{1} e^{x}dx\]

OpenStudy (jhannybean):

\[\large xe^{x}+e^{x} - e^{x}]_{0}^{1}\]

OpenStudy (jhannybean):

Input your values in the fashion \[\large F(b)-F(a)\] and you'll be set! :)

OpenStudy (anonymous):

ok Thank you!!

OpenStudy (jhannybean):

\[\large [(1)e^{(1)}+e^{1}-e^{1}]-[(0)e^{0}+e^{0}-e^{0}]\]

OpenStudy (jhannybean):

No problem :)